Oswal Mills Barnala consigned 5000 kg of vanaspati ghee to Rajendra
Dealers of Panipat. Each kg. Ghee costs Rs. 8. Oswal Mills paid Rs. 50 for
carriage, Rs. 250 for freight and Rs. 200 for insurance in transit. During
transit 500 kg. Ghee was accidentally destroyed for which insurance
company paid directly to the consignor Rs. 2500 in full settlement of the
claim. After 3 month from the date of consignment of goods to Panipat,
Rajendra Dealers reported that 1500 kg. Ghee was sold at Rs. 9.5 per kg.
The expenses were: On Godown Rent Rs. 500, On Salesman Salary Rs.
700. Rajendra dealers are entitled to a commission of 5% on sales. Due to
leakage, Rajendra Dealers also reported a loss of 20 kg. Ghee. Prepare
consignment account and abnormal loss account in the books of the
Consignor.
1
Expert's answer
2020-12-02T20:11:05-0500

OSWAL MILLS AND RAJENDRA PANIPAT DEALERS

CONSIGNMENT ACCOUNT

                                      Rs

 

Goods on consignment                         40,000

Carriage                                                     50

Freight charges                                          250

Insurance on transit                                   200

Rent                                                           500

Salesman salary                                         700

Commission 5% of 14250                         712.50

Profit                                                          323.50

 

TOTAL                                                     42,376

Rs

 

Abnormal loss                    4,050

Sales                                  14,250

Stock c/d                            24,436

 

 

 

 

 

 

TOTAL                                42,376

 

ABNORMAL LOSS ACCOUNT

Rs

 

To consignment account          4,050

 

 

4,050

                                            Rs

 

By insurance company                                  2,500

To profit and loss account transfer               1,550

 

   TOTAL                                           4,050

 

 

Workings

1. Abnormal loss has been calculated as follows

Cost of 5000kg @ RS 8…………………40,000

Freight…………………………………..250

Carriage…………………………………50

Insurance………………………………..200

TOTAL…………………………………40,500

Hence cost of the 500 kg wasted = (40,500/5000) ×500= Rs 4050

So the cost of goods involved in abnormal loss is Rs 4050

 

2. Value of closing stock has been computed as follows

Cost of 5000 kg @ RS 8…………………40,000

Freight…………………………………..250

Carriage…………………………………50

Insurance………………………………..200

Rent……………………………………..500

TOTAL cost of goods…………………..41,000

Cost per kg\ value of closing stock= 41,000/5,000 = Rs 8.2 ×2,980 kg= Rs 24,436.

Number of items of Closing stock

Total……5,000 kg

Less spoilt... (500 kg)

       Sold….(1500 kg)

       Leaked ...(20 kg)


                     2980 kg